### v1

# Data-Efficient Safe Policy Improvement Using Parametric Structure

Abstract
Safe policy improvement (SPI) is an offline reinforcement learning problem in which a new policy that reliably outperforms the behavior policy with high confidence needs to be computed using only a dataset and the behavior policy. Markov decision processes (MDPs) are the standard formalism for modeling environments in SPI. In many applications, additional information in the form of parametric dependencies between distributions in the transition dynamics is available. We make SPI more data-efficient by leveraging these dependencies through three contributions: (1) a parametric SPI algorithm that exploits known correlations between distributions to more accurately estimate the transition dynamics using the same amount of data; (2) a preprocessing technique that prunes redundant actions from the environment through a game-based abstraction; and (3) a more advanced preprocessing technique, based on satisfiability modulo theory (SMT) solving, that can identify more actions to prune. Empirical results and an ablation study show that our techniques increase the data efficiency of SPI by multiple orders of magnitude while maintaining the same reliability guarantees.
[…]
6125
1 Introduction
Reinforcement learning (RL) is the standard paradigm for solving sequential decision-making problems in unknown environments [42]. In an online RL setting, the agent interacts with an unknown environment to explore and gather information about its dynamics. This environment is typically modeled as a Markov decision process (MDP), and the agent’s goal is to compute a policy that maximizes the expected cumulative discounted reward [32].
Such online interaction is often undesirable or dangerous in real-world application areas such as robotics or healthcare [26, 21]. Online safe RL methods attempt to mitigate these concerns by constraining the agent’s interaction [17], for instance, through shielding [1] or constrained actor-critic methods [48]. Nevertheless, online safe RL approaches still fundamentally rely on the agent exploring its environment.
In settings where such exploration is infeasible, offline RL can serve as an alternative. Offline RL prevents the agent from interacting with the environment and instead only allows the use of a dataset of past interactions generated by some behavior policy [26, 22]. Using only this dataset, the agent is tasked with finding a policy that performs well. In settings where the behavior policy is also available, the task is to find a new policy that outperforms the behavior policy.
Safe policy improvement (SPI) [44, 29] is an offline RL problem that imposes a probabilistic correctness guarantee on the performance gain of the new policy over the behavior policy. Safe policy improvement with baseline bootstrapping (SPIBB) [24] is one of the most commonly used SPI methods. SPIBB ensures the improvement guarantee is satisfied by constraining the standard policy iteration algorithm to bootstrap to (i.e., follow) the behavior policy in state-action pairs where insufficient data is available. Note that most SPI methods that ensure the improvement guarantee rely on discrete environments [29, 24, 46, 10, 8, 37, 40].
Since the agent cannot collect any new data about the environment in SPI, it is paramount to make the most of the available information. Existing SPI methods scale poorly in terms of the number of samples required, as also evidenced by the practice of using hyperparameters instead of the theoretically required sample sizes, see, e.g.[24, 46]. This challenge to make SPI(BB) more data efficient has given rise to methods that exploit structure in the underlying MDP, such as factored state-spaces [38, 39] or graph connectivity [46].
While exact probabilities are often unknown, many real-world problems exhibit structure and dependencies in the underlying probability distributions. For instance, in robotics, probability distributions are used to represent many different sources of uncertainty [25]. One such source could be component failure, and the same component may fail with the same probability distribution across several different configurations, i.e., state-action pairs, and thus introduce dependencies between the probability distributions. For a concrete example, consider slippery grid worlds such as Frozen Lake, where a single slip probability affects multiple transitions [45].
Parametric MDPs (pMDPs) are an extension of MDPs suitable to model parametric dependencies between probability distributions [18, 3, 20]. In a pMDP, each transition is assigned a polynomial over parameters that can occur on multiple transitions, effectively defining dependencies between probability distributions on different state-action pairs.
Contributions
In this paper, we investigate how to improve the data efficiency of SPI methods using such parametric structures. We present three methods that exploit the additional knowledge given by the pMDP to make SPIBB more data efficient. Specifically, we introduce a parametric variant of SPIBB, as well as two preprocessing techniques that are independent of the dataset. In more detail, our contributions are:
Input Dataset Behavior Policy Parametric MDPGame-Based Pruning (Contribution 2)SMT-Based Pruning (Contribution 3)Existing SPI(BB) Methods, e.g., [24, 46]Parametric SPIBB (Contribution 1)Result Improved Policy
Figure 1: Overview of our contributions and how they relate.
  1. 1.
    Parametric SPIBB (Section 3). We extend SPIBB into parametric SPIBB (pSPIBB) that exploits the parametric structure of the given pMDP. Specifically, when probability distributions at different state-action pairs follow the same parametric structure, we may lump together samples from these state-action pairs. This effectively increases the available data, and as a result, pSPIBB has fewer state-action pairs that need to be bootstrapped to the behavior policy compared to standard SPIBB.
  2. 2.
    Game-based pruning (Section 4). Our second contribution is a preprocessing method that prunes state-action pairs from the environment via a game-based abstraction. By reasoning over all possible probability distributions, we obtain lower and upper bounds on the value of the pMDP. Using these bounds, we prune sub-optimal state-action pairs for which we can guarantee that an optimal policy will never visit these state-action pairs.
  3. 3.
    SMT-based pruning (Section 5). Our third contribution is an exact pruning method based on a satisfiability modulo theory (SMT) query. Since game-based pruning is an abstraction technique, it can be overly conservative. In contrast, SMT-based pruning reasons directly over the transition polynomials, and thus potentially removes more state-action pairs than game-based pruning, but at additional computational expense.
Figure 1 provides an overview of how our contributions relate. In particular, standard SPI(BB) methods can be used on the input by simply ignoring the parametric MDP, while pSPIBB takes the parametric structure into account. Additionally, the two pruning methods serve as preprocessing steps for both standard SPI(BB) methods and pSPIBB. In Section 6, we provide an extensive empirical evaluation and ablation study for these techniques on four well-known benchmarks and a newly introduced Rock-Paper-Scissors environment. All proofs can be found in the appendix.
2 Background
By $|X|$, we denote the number of elements in a finite set $X$. A discrete probability distribution over a finite set $X$ is a function $\mu\colon X\to[0,1]$ with $\sum_{x\in X}\mu(x)=1$. The set of all distributions over $X$ is denoted $\Delta(X)$. The support of a probability distribution $\mu$ is the set of all elements with non-zero probability: $\mathrm{supp}(\mu)=\{x\in X\mid\mu(x)>0\}$.
Definition 1 (MDP).
A Markov decision process (MDP) is a tuple $\langle S,\iota,A,T,R,\gamma\rangle$, where $S$ is a finite set of states, $\iota\in S$ is the initial state, $A$ is a finite set of actions, $T\colon S\times A\to\Delta(S)$ is the probabilistic transition function, $R\colon S\times A\to[-R_{\max},R_{\max}]$ is the (bounded) reward function, and $\gamma\in(0,1)$ is the discount factor.
We write $T(s^{\prime}\!\mid\!s,a)$ for the probability $T(s,a)(s^{\prime})$. By $\mathrm{supp}(s,a)=\left\{s^{\prime}\!\mid\!T(s^{\prime}\!\mid\!s,a)>0\right\}$ we denote the support of a state-action pair $\langle s,a\rangle$. A trajectory is a sequence of successive states and actions: $s_{0}a_{0}s_{1}\dots$, such that $s_{i+1}\in\mathrm{supp}(s_{i},a_{i})$ and $s_{0}=\iota$.
A stationary (or memoryless) stochastic policy is a function $\pi\colon S\to\Delta(A)$. The objective is to find a policy that maximizes the expected cumulative discounted reward:
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right],$
where $\langle s_{t},a_{t}\rangle$ is the state-action pair visited at time $t$ when following $\pi$. The value of an MDP is the unique least fixed point of the following recursive equations:
$Q(s,a)$ $=R(s,a)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}\!\mid\!s,a)V(s),$ (1)
$V(s)$ $=\max_{a\in A}Q(s,a),$ (2)
where $Q(s,a)$ are called the state-action values and $V(s)$ are the state-values, both initialized at zero. We write $V^{\pi}$ and $Q^{\pi}$ for both value functions under some policy $\pi$, which we call the performance of $\pi$. Optimal state and state-action values are given by the performance of an optimal policy $\pi^{*}$, denoted by $V^{*}=V^{\pi^{*}}$ and $Q^{*}=Q^{\pi^{*}}$, respectively.
2.1 Safe Policy Improvement
The offline RL problem of safe policy improvement (SPI) is to compute a policy $\pi_{I}$ that improves over a given behavior policy $\pi_{B}$, based on historical data $\mathcal{D}=\braket{s_{i},a_{i},s_{i+1}}_{i=0\dots n}$ obtained from executing $\pi_{B}$. Specifically, for an admissible performance loss $\zeta$ and confidence parameter $\delta$, the performance of the improved policy $\pi_{I}$ must be better than that of $\pi_{B}$ with high probability:
$\operatorname{\mathrm{Pr}}\big{(}V^{\pi_{I}}(\iota)\geq V^{\pi_{B}}(\iota)-\zeta\big{)}>1-\delta.$ (3)
For a dataset $\mathcal{D}$, we denote by #$\#_{\mathcal{D}}(s,a)$ and #$\#_{\mathcal{D}}(s,a,s^{\prime})$ the number of times a state-action pair or transition occurs in $\mathcal{D}$. The maximum likelihood estimate MDP (MLE-MDP) of a dataset $\mathcal{D}$ is given by $\tilde{M}=\langle S,\iota,A,\tilde{T},R,\gamma\rangle$, where
$\tilde{T}(s^{\prime}\!\mid\!s,a)=\frac{\#_{\mathcal{D}}(s,a,s^{\prime})}{\#_{\mathcal{D}}(s,a)}$ (4)
when #$\#_{\mathcal{D}}(s,a)>0$, and zero otherwise. [^1] Note that when no data is available for a state-action pair, that action is disabled in the MLE-MDP.
The safe policy improvement with baseline bootstrapping (SPIBB) [24] algorithm solves the SPI problem by only deviating from the behavior policy in states where sufficient data is available. Specifically, SPIBB defines a set $\mathcal{U}$ of state-action pairs that are uncertain. That is, state-action pairs of which the number of samples contained in $\mathcal{D}$ is below a specified threshold $N_{\wedge}$:
$\mathcal{U}=\left\{\braket{s,a}\in S\times A\mid\#_{\mathcal{D}}(s,a)<N_{\wedge}\right\}.$ (5)
The improved policy $\pi_{I}$ is computed as a policy that chooses an action with the probability of $\pi_{B}$ if that state-action pair is uncertain, and concentrates the remaining probability mass on an action that is optimal for the MLE-MDP. Formally, let $Q^{*}_{\tilde{M}}$ be the optimal state-action values of the MLE-MDP and let $\bar{A}(s)=\{a\in A\!\mid\!\langle s,a\rangle\not\in\mathcal{U}\}$ denote the non-bootstrapped actions at state $s$, the improved policy $\pi_{I}$ is defined as
$\pi_{I}(a\!\mid\!s)=$$\begin{cases}\pi_{B}(a\!\mid\!s)&\forall\langle s,a\rangle\in\mathcal{U},\\ \sum\limits_{a^{\prime}\in\bar{A}(s)}\pi_{B}(a^{\prime}\!\mid\!s)&a=\operatorname*{arg\,max}\limits_{a^{\prime}\in\bar{A}(s)}Q^{*}_{\tilde{M}}(s,a^{\prime}),\\ 0&\text{ otherwise}.\end{cases}$
For a fixed $\zeta$ and $\delta$, the threshold $N_{\wedge}$ required to ensure the SPI guarantee (Equation 3) can be computed by a binary search [46] and is bounded by
[…]
where $V_{\textit{max}}\leq\nicefrac{{R_{\max}}}{{1-\gamma}}$ is a bound on the maximal value. In practice, $N_{\wedge}$ is often set as a hyperparameter, together with $\delta$, and the possible loss $\zeta$ is computed instead [24]. For a fixed $N_{\wedge}$ and $\delta$[46] establish the currently best-known bound on $\zeta$ through the inverse incomplete Beta function [43] $I^{-1}$, being:
[…]
where $\delta_{T}=\nicefrac{{\delta}}{{2|S|^{2}|A|^{2}}}$ and $c=V_{\tilde{M}}^{\pi_{B}}(\iota)-V_{\tilde{M}}^{\pi_{I}}(\iota)$ is a constant for the difference in performance between the behavior policy $\pi_{B}$ and the improved policy $\pi_{I}$ on the MLE-MDP $\tilde{M}$.
2.2 Parametric MDPs
Parametric MDPs extend ordinary MDPs by a parametric transition function that maps transitions $\langle s,a,s^{\prime}\rangle$ to polynomials with rational coefficients over a set of variables $X$. We write $\mathbb{Q}[X]$ for the set of all such polynomials.
Definition 2 (pMDP).
A parametric MDP (pMDP) is a tuple $\langle S,\iota,A,P,X,R,\gamma\rangle$, where $S$, $\iota$, $A$, $R$, and $\gamma$ are as in MDPs, $X$ is a set of parameters and $P\colon S\times A\times S\to\mathbb{Q}[X]$ is the parametric transition function.
We assume that all transition polynomials that are not (syntactically) 00 have some nonzero value in their image. We write $\mathrm{supp}(s,a)=\left\{s^{\prime}\!\mid\!P(s,a,s^{\prime})\neq 0\right\}$. A pMDP can be instantiated into an MDP by applying a valuation $\theta\in\mathbb{R}^{\lvert X\rvert}$, which results in an ordinary transition function $P_{\theta}(s^{\prime}\!\mid\!s,a)=P(s,a,s^{\prime})(\theta)$. We extend state and state-action values to pMDPs instantiated by a valuation $\theta$, denoted $V_{\theta}$ and $Q_{\theta}$, respectively, as well as $Q^{\pi}_{\theta}$ and $V^{\pi}_{\theta}$ for the performance of a policy $\pi$. We only consider graph-preserving valuations $\theta\in\Theta_{\mathrm{gp}}$, i.e., valuations that all share the same underlying graph. Formally, all $\theta\in\Theta_{\mathrm{gp}}$ conform to the following, for all $s,s^{\prime}\in S$ and all $a\in A$: (1) $\sum_{s^{\prime\prime}\in S}P_{\theta}(s,a,s^{\prime\prime})=1$, (2) $0\leq P_{\theta}(s,a,s^{\prime})\leq 1$, and (3) $P(s,a,s^{\prime})\neq 0\implies P_{\theta}(s,a,s^{\prime})\neq 0$.
3 Parametric SPIBB
As the introduction outlines, our first contribution is the extension of SPIBB into parametric SPIBB (pSPIBB), which relies on the following two modifications. First, we adapt the MLE-MDP (Equation 4) to consider occurrences of polynomials instead of transitions in the dataset. Second, we replace the set of uncertain state-action pairs (Equation 5) by again counting the occurrences of polynomials instead of state-action pairs. We discuss these modifications in detail in the remainder of this section.
We assume that state-action distributions do not label transitions with identical polynomials, i.e., $P(s,a,s^{\prime})=P(s,a,s^{\prime\prime})$ implies $s^{\prime}=s^{\prime\prime}$. This assumption is without loss of generality as successors along a transition with identical labels can be replaced by a fresh intermediate state followed by a uniform distribution over the original successors (cf. reduction to two-successor MDPs [46]). We lift the polynomial labeling to state-action pairs by setting $P(s,a)=\{P(s,a,s^{\prime})\mid s^{\prime}\in S\}$.
We define equivalence classes of state-action pairs and transitions and treat the polynomials as labels to count their occurrences together in the data, a technique known as parameter tying [36, 31, 30]. Note that we do not need to know the exact transition probabilities, since we are only comparing the polynomial labels of the given pMDP. We write $[s,a]_{P}$ for the set $\{\langle q,b\rangle\!\mid\!P(s,a)=P(q,b)\}$ and $[s,a,s^{\prime}]_{P}$ for $\{\langle q,b,q^{\prime}\rangle\!\mid\!P(s,a,s^{\prime})=P(q,b,q^{\prime})\text{ and }P(s,a)=P(q,b)\}$. Note that equivalence between transitions also requires the state-action pair to be equivalent in terms of polynomial labels over successor transitions (see Figure 2). We define the MLE-MDP over a pMDP by replacing Equation 4 by:
$\tilde{P}(s^{\prime}\!\mid\!s,a)=\frac{\sum_{(q,b,q^{\prime})\in[s,a,s^{\prime}]_{P}}\#_{\mathcal{D}}(q,b,q^{\prime})}{\sum_{(q,b)\in[s,a]_{P}}\#_{\mathcal{D}}(q,b)}.$ (6)
$s_{0}$$s_{1}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$$s_{6}$$s_{7}$$a$$b$$c$$x$$1-x$$x$$y$$1-x-y$$y$$1-y$
Figure 2: An example pMDP with parameters $x$ and $y$. In this pMDP, pSPIBB’s MLE-MDP does not reduce uncertainty, e.g., $P(s_{0},a,s_{1})=P(s_{0},b,s_{3})$ but $\langle s_{0},b,s_{3}\rangle\not\in[s_{0},a,s_{1}]_{P}$ because $P(s_{0},b,s_{5})\not\in[s_{0},a]_{P}$. Further note that if #$\nicefrac{{\#_{\mathcal{D}}(s_{0},c,s_{6})}}{{\#_{\mathcal{D}}(s_{0},c)}}=\nicefrac{{\#_{\mathcal{D}}(s_{0},a,s_{1})}}{{\#_{\mathcal{D}}(s_{0},a)}}=1$, it is unclear what the values of $\tilde{P}(s_{i}\!\mid\!s_{0},b)$ should be set to in order to minimize uncertainty while obtaining a valid distribution.
We modify SPIBB into pSPIBB by taking shared parameters into account. We first replace Equation 4 with Equation 6. Then we redefine the uncertainty set by taking into account not only how many times a state-action pair $\langle s,a\rangle$ occurs in the dataset but also how many identically-labeled pairs $\langle q,b\rangle\in[s,a]_{L}$ do:
$\mathcal{U}=\Big{\{}\braket{s,a}\in S\times A\>\Big{|}\>\sum_{(q,b)\in[s,a]_{L}}\#_{\mathcal{D}}(q,b)<N_{\wedge}\Big{\}}.$ (7)
As a result, this new uncertainty set is a subset of the original uncertainty set that does not use parameter sharing (Equation 5); hence, pSPIBB is guaranteed to perform equally well or better compared to standard SPIBB.
4 Game-Based Analysis
Our second contribution makes both SPIBB and pSPIBB more data efficient by pruning state-action pairs for which we can guarantee that an optimal policy will never visit these state-action pairs. First, we analyze sequential decision-making in (p)MDPs from a game-based perspective and then show how to use a game-based abstraction scheme to prune unnecessary state-action pairs.
Decision-making under uncertainty can be phrased as a two-player game between an agent and nature [28]. The agent selects the actions, while nature draws successor states from the probability distributions in a (p)MDP. Taking this two-player game point of view, we can replace nature with cooperative and antagonistic players to obtain upper and lower bounds on the values that can be achieved in the unknown pMDP. Intuitively, playing against a cooperative player will result in an optimistic overapproximation of the value, while playing against an antagonistic player yields a worst-case underapproximation. These bounds hold regardless of the (unknown) transition probabilities that influence the behavior of nature. Hence, we can use the bound to prune the set of uncertain state-action pairs in SPI settings, even in the complete absence of data.
Playing against a cooperative or antagonistic nature, the agent’s respective optimal policies are defined as:
$\overline{\pi}={}$ $\operatorname*{arg\,max}_{\pi}\Big{(}\sup_{\rho}\sum\nolimits_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\Big{)},$
$\underline{\pi}={}$ $\operatorname*{arg\,max}_{\pi}\Big{(}\inf_{\rho}\sum\nolimits_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\Big{)},$
where the supremum and infinum range over trajectories $\rho=s_{0}a_{0}s_{1}\dots$ consistent with $\pi$, i.e., $a_{i}\in\mathrm{supp}(\pi(s_{i}))$ for all $i\geq 0$, and $\langle s_{t},a_{t}\rangle$ is the state-action pair visited at time $t$ along the run. We now define the antagonistic and cooperative values of a pMDP.
Definition 3 (aVal and cVal).
The antagonistic and cooperative values of the pMDP are the unique solutions to the corresponding Bellman equations [49, Theorem 5.1]:
$\mathbf{aVal}(s)$ $=\max_{a\in A}\min_{s^{\prime}\in\mathrm{supp}(s,a)}R(s,a)+\gamma\mathbf{aVal}(s^{\prime}),$ (8)
$\mathbf{cVal}(s)$ $=\max_{a\in A}\max_{s^{\prime}\in\mathrm{supp}(s,a)}R(s,a)+\gamma\mathbf{cVal}(s^{\prime}).$ (9)
We omit $\theta$ from the notation since the values depend only on the support of the probability distributions and not on the actual probabilities themselves. We write $\mathbf{aVal}^{\pi}$ and $\mathbf{cVal}^{\pi}$ for the antagonistic and cooperative value functions, respectively, under policy $\pi$, i.e., the unique solution to:
$\mathbf{aVal}^{\pi}(s)$ $=\min_{\begin{subarray}{c}a\in\mathrm{supp}(\pi(s))\\ s^{\prime}\in\mathrm{supp}(s,a)\end{subarray}}R(s,a)+\gamma\mathbf{aVal}^{\pi}(s^{\prime}),$ (10)
$\mathbf{cVal}^{\pi}(s)$ $=\max_{\begin{subarray}{c}a\in\mathrm{supp}(\pi(s))\\ s^{\prime}\in\mathrm{supp}(s,a)\end{subarray}}R(s,a)+\gamma\mathbf{cVal}^{\pi}(s^{\prime}).$ (11)
Note that $\mathbf{aVal}^{\underline{\pi}}(\iota)=\mathbf{aVal}(\iota)$ and $\mathbf{cVal}^{\overline{\pi}}(\iota)=\mathbf{cVal}(\iota)$.
For intuition with respect to Equation 10 and Equation 11, note that under policy $\pi$, all remaining probabilistic choices are resolved antagonistically or cooperatively, respectively. This resolution can be seen as modifying the policy to choose the worst, respectively, best, actions only—instead of having a distribution over multiple actions.
We use these antagonistic and cooperative values to find sub-optimal state-action pairs, i.e., pairs $\langle s,a\rangle$ for which we can guarantee that an optimal policy $\pi^{*}$ will never select $a$ in state $s$. In that case, we can prune $\langle s,a\rangle$ from the pMDP.
To establish the correctness of this approach, we first present two results that describe the relation between the expected discounted reward and the antagonistic values of a state. The first result, Lemma 1, states that $\mathbf{aVal}$ and $\mathbf{cVal}$ are inclusive lower and upper bounds on the expected discounted reward $V_{\theta}$.
Lemma 1.
Given a pMDP, the following inequalities hold for all $s\in S$ and all graph-preserving valuations $\theta\in\Theta_{\mathrm{gp}}$.
$\mathbf{aVal}(s)\leq V_{\theta}(s)\leq\mathbf{cVal}(s).$
The proof immediately follows from the definitions and properties of convex combinations.
The second result, Theorem 1, provides an effective characterization of when a policy almost-surely, i.e., with probability one, obtains a cumulative reward that is strictly higher than the antagonistic value.
Let =$\mathcal{I}=\{\langle s,a,s^{\prime}\rangle\in S\times A\times S\!\mid\!\mathbf{aVal}(s)<R(s,a)+\gamma\mathbf{aVal}(s^{\prime})\}$ denote the set of transitions along which the antagonistic value is strictly improved. We write $\tau_{\mathcal{I}}$ for the first hit time of $\mathcal{I}$, i.e., the random variable for the first time step where a transition from $\mathcal{I}$ is traversed. The first hit time is $\infty$ if such a transition is never traversed. Intuitively, we show that a policy almost-surely obtains a cumulative discounted reward that is strictly larger than the antagonistic value if and only if it is a worst-case optimal policy and it almost-surely improves the antagonistic value along some transition.
Theorem 1.
Let $\theta\in\Theta_{\mathsf{gp}}$. Then, there is a policy $\pi$ such that
$\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\sum\nolimits_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})>\mathbf{aVal}(\iota)\right)=1,$ (12)
if and only if there is a worst-case optimal policy $\pi$, i.e., $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$, with $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\right)=1$.
Almost-sure finite-time hitting (also known as reachability) does not depend on concrete probability values but only on their support [2, 7]. Since the antagonistic value is independent of $\theta$ (and can be computed, e.g., as a solution to Equation 8), we get that Equation 12 is also independent of $\theta$ and can be checked on the given pMDP for a policy $\pi$.
As an immediate consequence of Theorem 1, we get that the lower bound from Lemma 1 is strict under some conditions that are easy to check, as stated in the following lemma.
Lemma 2.
If there exist $\theta\in\Theta_{\mathrm{gp}}$ and a policy $\pi$ such that $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\right)=1$ then for all $\theta\in\Theta_{\mathrm{gp}}$ we have $\mathbf{aVal}(\iota)<V_{\theta}(\iota)$.
It is easy to check, using Equation 8, that any policy satisfying $\mathbf{aVal}(s)=\min_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{aVal}(s^{\prime})$ for all $s$ and all $a\in\mathrm{supp}(\pi(s))$ is worst-case optimal, i.e., $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$. Conversely, for any worst-case optimal policy $\pi$, we can make local modifications to obtain a second worst-case optimal policy $\pi^{\prime}$ satisfying the equations. These observations suggest the following two-step procedure to determine whether the conditions of Lemma 2 hold:
  1. 1.
    Remove all state-action pairs $\langle s,a\rangle$ for which we have $\mathbf{aVal}(s)>\min_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{aVal}(s^{\prime})$;
  2. 2.
    In the resulting pMDP, use Algorithm 45 for almost-sure reachability from [2] to determine whether a transition from $\mathcal{I}$ is hit with probability $1$ under some policy $\pi$.
The first step yields a sub-pMDP in which all policies are worst-case optimal. Moreover, any worst-case optimal policy has a corresponding policy in this sub-MDP. The second step tells us whether a policy to hit the improving transition almost-surely exists. Since the second step is realized on the sub-pMDP obtained from step $1$, we get a positive answer if and only if the conditions from Lemma 2 hold true (see Appendix D).
4.1 Game-Based Pruning
We use our results on game values to reduce the number of state-action pairs in the pMDP. Leveraging the bounds from Lemmas 1 and 2, we compare the antagonistic value of a state with the cooperative values of its successor states.
Theorem 2 (aVal-cVal pruning).
Let $\pi$ be a policy and $\langle s,a\rangle$ be a state-action pair. If $a\in\mathrm{supp}(\pi(s))$ and
$\mathbf{aVal}(s)>\max_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{cVal}(s^{\prime}),$ (13)
then for all $\theta\in\Theta_{\mathrm{gp}}$ we have $V_{\theta}(s)>V_{\theta}^{\pi}(s)$. Moreover, if $\mathbf{aVal}^{\pi}(s)=\mathbf{aVal}(s)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\!\mid\!s_{0}=s\right)=1$, for some $\pi$, the non-strict version of Equation 13 suffices.
Essentially, Theorem 2 states that the action $a$ will not be chosen in state $s$ by any optimal policy and can therefore be pruned. Formally, if the condition of the result is true then, by Equation 9, any policy $\pi$ that plays $a$ from $s$ has a performance $Q_{\theta}^{\pi}(s,a)$ that is strictly smaller than the (optimal) antagonistic value $\mathbf{aVal}(s)$. By Lemma 1, it then follows that its performance $V^{\pi}_{\theta}(s)$ is suboptimal, irrespective of the valuation $\theta\in\Theta_{\mathrm{gp}}$. The second part follows from Lemma 2.
$s_{1}$$s_{0}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$$s_{6}$$s_{7}$$s_{8}$$s_{9}$$a,r{=}0$$b,r{=}0$$c,r{=}0$$r{=}0$$r{=}5$$r{=}0$$r{=}0$$r{=}{-}1$$r{=}{-}5$$r{=}{-}10$$r{=}{-}10$$r{=}0$
Figure 3: Example of a pMDP where both cases of aVal-cVal pruning apply (see Theorem 2).
Example 1.
Consider the pMDP in Figure 3 and let $\gamma=0.95$. First, note that $\mathbf{aVal}(s_{0})=\mathbf{aVal}(s_{1})=\mathbf{cVal}(s_{3})=\mathbf{cVal}(s_{4})=0$ and $\mathbf{cVal}(s_{7})=\mathbf{cVal}(s_{8})=-200$. Therefore, $\mathbf{cVal}(s_{5})=-191$ and $\mathbf{cVal}(s_{6})=-195$. We thus have $\mathbf{aVal}(s_{0})>-1+\gamma\mathbf{cVal}(s_{5})$ and $\mathbf{aVal}(s_{0})>-5+\gamma\mathbf{cVal}(s_{6})$. Therefore, by Theorem 2, the state-action pair $\langle s_{0},c\rangle$ can be removed. Second, we have that the antagonistic value strictly improves along $\langle s_{2},a,s_{9}\rangle$, i.e., it is an element of $\mathcal{I}$, and it can be hit almost surely by playing action $a$ from $s_{0}$. Therefore, the condition from the second part of Theorem 2 is satisfied and $\langle s_{0},b\rangle$ can also be removed, since $\mathbf{aVal}(s_{0})=0+\gamma\mathbf{cVal}(s_{3})$ and $\mathbf{aVal}(s_{0})>0+\gamma\mathbf{cVal}(s_{4})$.
5 Parametric Value Function Analysis
In all techniques presented thus far, we have not yet fully exploited the parametric structure of pMDPs. Even in our definition of an MLE-MDP for pSPIBB, Equation 6, we do not yet consider all constraints implied by the sharing of parameters across transitions.
We proceed in this direction by leveraging the fact that both the state and state-action value functions $V_{\theta}$ and $Q_{\theta}$ are known to be rational functions of $\theta$ [14, 23]. For intuition, observe that the systems from Equations 2 and 1 can be solved via algebraic manipulations, noting that probabilities are now polynomials over $\theta$. A complete algorithm to compute both functions is given in [3, Section 3].
Instead of computing the rational functions, we encode the systems of polynomial constraints from Equations 2 and 1 and use a satisfiability modulo theory (SMT) solver to determine the (im)possibility of uncertain state-action pairs being useful to obtain $\pi_{I}$. In a nutshell, SMT solvers can check the satisfiability of first-order formulas over some theory [5]. Here, the relevant theory is the existential theory of the reals (ETR), which allows rational constants, multiplication, addition, and existentially quantified real-valued variables, see [35] for decision problems that can be encoded in ETR, including the encoding for pMDPs that we are alluding to.
5.1 SMT-Based Pruning
We now present a version of Theorem 2 that compares the maximal value of a state over all possible valuations $\theta\in\Theta_{\mathrm{gp}}$ against its antagonistic value. Instead of the cooperative value, we now use a (sometimes) tighter upper bound on the value of the state to compare against the antagonistic value.
Theorem 3 (aVal-Q pruning).
Let $\pi$ be a policy, $\theta\in\Theta_{\mathrm{gp}}$, and $\langle s,a\rangle$ a state-action pair. If $a\in\mathrm{supp}(\pi(s))$ and
$\mathbf{aVal}(s)>Q_{\theta}(s,a),$ (14)
then $V_{\theta}(s)>V_{\theta}^{\pi}(s)$. Moreover, if $\mathbf{aVal}^{\pi}(s)=\mathbf{aVal}(s)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\!\mid\!s_{0}=s\right)=1$, for some $\pi$, the non-strict version of Equation 14 suffices.
The argument why the result holds is similar to the one for Theorem 2. If the condition holds, by (the parametric versions of) Equations 1 and 2, all policies $\pi$ that play $a$ from $s$ have a performance $V_{\theta}^{\pi}(s)$ strictly smaller than $\mathbf{aVal}(s)$. Thus, by Lemma 1, its performance $V^{\pi}_{\theta}(s)$ is suboptimal. The complement of Equation 14, with valuations $\theta$ quantified existentially and constraints on $\theta$ to be an element of $\Theta_{\mathrm{gp}}$, can be encoded in the existential theory of the reals.
Finally, observe that Theorem 3 can be further improved by comparing the values of state-action pairs against each other instead of against the antagonistic value of their source state. Instead of precomputing $\mathbf{aVal}(s)$, we can replace the left-hand side of Equation 14 by the (encoding of) the actual value of the state. Correctness of the result follows from (the parametric versions of) Equations 1 and 2 directly.
Theorem 4 (Q-Q pruning).
Let $\pi$ be a policy, $\theta\in\Theta_{\mathrm{gp}}$, and $\langle s,a\rangle$ a state-action pair. If $a\in\mathrm{supp}(\pi(s))$ and $V_{\theta}(s)>Q_{\theta}(s,a)$, then $V_{\theta}(s)>V_{\theta}^{\pi}(s)$.
$s_{0}$$s_{1}$$s_{2}$$s_{3}$$a$$b$$a$$p$$1-p$$p$$1-p$$1-p$$p$$r{=}20$$r{=}{-}5$
Figure 4: Example of a pMDP where aVal-Q pruning fails but Q-Q pruning does apply (see Theorem 3 and Theorem 4).
Example 2.
Consider the pMDP in Figure 4 and let $\gamma=0.9$. Worst case, state $s_{3}$ is reached after playing the first action. Hence, $\mathbf{aVal}(s_{0})=0.9(-5)(1-0.9)^{-1}=-45<Q_{\theta}(s_{0},a),Q_{\theta}(s_{0},b)$, for all $\theta\in\Theta_{\mathrm{gp}}$, since playing either action yields some positive probability of reaching state $s_{2}$, where the expected discounted reward is positive. Hence, we cannot use Theorem 3 to remove any action. By the same reasoning, however, we have $Q_{\theta}(s_{0},a)\geq p^{2}$ and $p\geq Q_{\theta}(s_{0},b)$. Since $p>p^{2}$ for all graph-preserving valuations, i.e., for all $p\in(0,1)$, then $V_{\theta}(s_{0})=Q_{\theta}(s_{0},b)<Q_{\theta}(s_{1},a)$, thus by Theorem 4, we can remove $\langle s_{0},a\rangle$.
(a) Taxi, $N_{\wedge}=200$.
(b) Gridworld, $N_{\wedge}=200$.
(c) Resource gathering, $N_{\wedge}=200$.
(d) Pac-Man, $N_{\wedge}=200$.
(e) Rock-paper-scissors, $N_{\wedge}=200$.
(f) Rock-Paper-Scissors, $N_{\wedge}=1000$.
Figure 5: Comparison of methods on different environments for a specified $N_{\wedge}$ value.
6 Experimental Evaluation
We empirically evaluate our contributions, pSPIBB, game-based pruning, and SMT-based pruning, compared to standard SPIBB [24].
Environments.
We use the following standard benchmarks in our evaluation: Resource Gathering [6], Gridworld [24], Pac-Man [19], and Taxi [16]. Additionally, we implement a biased version of the Rock-Paper-Scissors game as a new benchmark. Full details on these environments can be found in Appendix H.
6.1 Feasibility of SMT-Based Pruning
We start with evaluating the feasibility of SMT-based pruning. Since SMT solving is known to be computationally expensive, we pose the research question: RQ1: Is SMT-based pruning practically feasible?
Setup.
We run SMT-based pruning on Gridworld to establish whether a single state-action pair can be pruned, using the SMT-solvers Z3 [15] and cvc5 [4]. We set a timeout of $12$ hours. Both Z3 and cvc5 were invoked via their Python interfaces with default options. We did not use any advanced encodings or solver settings.
Results and Discussion.
After a computation time of $6$ hours, Z3 establishes that the state-action pair cannot be pruned, while cvc5 timed out after $12$ hours. From these results, we conclude that SMT-based pruning is (at present) practically infeasible, answering RQ1 negatively. We emphasize, however, that our encoding of the conditions of Theorems 3 and 4 is quadratic in the number of states, actions, and transitions of the pMDP, and may be improved. Alternatively, tailor-made optimizations for these specific SMT instances could exist.
6.2 Ablation of pSPIBB and Game-Based Pruning
Next, we evaluate and compare the effects of our contributions regarding data efficiency. Since we already established that SMT-based pruning is not practically feasible in Section 6.1, we omit this method from the comparison. The central research question of this ablation study is RQ2: Are pSPIBB, game-based pruning, and their combined use more data efficient than standard SPIBB?
Setup.
We generate datasets of increasing size for each environment and compute the improved policy’s performance across those datasets using each (combination of) our methods and standard SPIBB. Whenever pSPIBB or game-based pruning has no (additional) effect because there are no parameters shared or state-action pairs to prune, we omit those curves from the plots. For all environments, we repeat the experiment for $1024$ different seeds. We report the mean performance as well as the $10\%$ and $1\%$ conditional value at risk (CVaR) [33], which is the mean of the $10\%$ (resp. $1\%$) worst performing seeds.
6.2.1 Results and Discussion
Figure 5 shows the results obtained on all environments for $N_{\wedge}=200$, with an additional experiment using $N_{\wedge}=1000$ for Rock-Paper-Scissors. The Taxi environment is the only benchmark that contains shared parameters and where game-based pruning removes state-action pairs. There are no shared parameters in Pac-Man, and game-based pruning does not affect the other three environments.
Pruning and pSPIBB increase data efficiency.
In Figures 5(a) and 5(d) we see how pruning is an effective preprocessing method that can improve SPIBB’s data efficiency by two orders of magnitude. In the other environments, pruning had no effect. Exploiting shared parameters as done in pSPIBB also increases data efficiency compared to standard SPIBB, as seen in Figures 5(a), 5(e), 5(f), 5(b) and 5(c). The Pac-Man environment does not contain any shared parameters. Thus, pSPIBB coincides with SPIBB in this environment. Finally, as showcased in Figure 5(a), the combination of pSPIBB with game-based pruning delivered the greatest improvement in data efficiency, highlighting the benefits of combined usage of our contributions and positively answering RQ2.
Discussion of decreased performance.
In Taxi, Resource Gathering, and Rock-Paper-Scissors, a decrease in performance compared to the behavior policy occurs for some seeds, as evidenced by the $10\%$ and $1\%$-CVaR curves. These decreases are more severe when using pSPIBB compared to standard SPIBB. This behavior occurs because a wrong maximum likelihood estimate based on the data affects multiple state-action pairs simultaneously, exacerbating effects that may also occur in standard SPIBB. Nonetheless, we emphasize the following two observations. First, increasing the $N_{\wedge}$ hyperparameter suppresses this effect while still maintaining the gains in data efficiency of our contributions, as evidenced by Figures 5(e) and 5(f), and second, all observed decreases in performance are still within the allowed performance loss $\zeta$. For instance, assuming an error $\delta=0.05$, we obtain in the Taxi environment an allowed loss $\zeta\approx 2895.02$ and in Rock-Paper-Scissors we obtain $\zeta\approx 41.36$ and $\zeta\approx 19.09$ for $N_{\wedge}=200$ and $1000$, respectively. All curves are well within the range of the behavior policy’s performance ±$\pm\zeta$.
7 Related Work
In addition to the main SPI works mentioned in the introduction, significant work has been done to make SPI methods applicable in broader settings. Most notably, multi-objective [34], non-stationary [11], partially observable [40], and multi-agent [8] settings. For a recent overview of SPI methods, we refer to [37].
Since collecting more data is not an option, data efficiency is of great importance in offline RL settings such as SPI. To that end, (offline) RL methods have been integrated with additional domain knowledge such as factored state spaces [38] or (parametric) graph structures [39, 36], similar to our approach. Alternatively, less conservative statistical methods to derive guarantees have been used in SPI [46] and more general learning settings [27]. Another bottleneck of SPI(BB) is that most methods are tabular, i.e., they explicitly build an MLE-MDP from the dataset. To alleviate this drawback, [10] employ Monte Carlo Tree Search to enable SPIBB in large state spaces.
Game-based abstractions have been used in safety-constrained planning for partially observable MDPs [12]. This work was later extended to probabilistic constraints in [13].
8 Conclusion
We presented three new contributions that exploit parametric dependencies in the underlying environment to make safe policy improvement more data-efficient. Our experimental evaluation showed that parametric SPIBB and game-based pruning (and their combination) are practical and efficient techniques, increasing data efficiency by two orders of magnitude. In contrast, SMT-based pruning, while theoretically promising, is practically infeasible at present.
Future work will focus on increasing the computational efficiency of SMT-based pruning using more advanced encodings or other solvers. In this direction, we highlight that our code is agnostic with respect to the actual SMT-solver used. Integrating additional SMT-solvers can be done by simply implementing the small SMT-solver interface we provided. New benchmarks can also be easily added. Additionally, we aim to further increase the data-efficiency of SPI(BB) by integrating our methods with others such as structure learning [39]. Another direction is to overapproximate the parametric distributions by intervals, reducing it to a robust MDP [47, 41]. Such an approximation would increase computational efficiency at the cost of being more conservative, where the exact benefits of this trade-off remain an open question.
{ack}
This work was supported by the FWO “SynthEx” project (G0AH524N). We would also like to thank Nils Jansen and Sebastian Junges for their feedback on earlier versions of this work.
Appendix A Proof of Lemma 1
Consider a pMDP $\langle S,\iota,A,P,X,R,\gamma\rangle$. Let $s\in S$ and $\theta\in\Theta_{\mathrm{gp}}$ be arbitrary. We will now prove that the following inequalities hold:
$\mathbf{aVal}(s)\leq V_{\theta}(s)\leq\mathbf{cVal}(s).$
First, recall that:
$V_{\theta}(s)=\sup_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right],$
where the expectation is taken over the usual probability space, where trajectories are outcomes for Markov chains (indeed, having fixed $\pi$ in the MDP induced by the valuation $\theta$, we are dealing with a Markov chain). Further, recall that the expectation operator is monotonic. That is, for arbitrary random variables $X$ and $Y$, if $X\leq Y$ holds with probability $1$ and $\mathbb{E}[X],\mathbb{E}[Y]$ both exist then $\mathbb{E}[X]\leq\mathbb{E}[Y]$.
Now, let $\pi$ be an arbitrary policy and $\rho=s_{0}a_{0}s_{1}$ a trajectory consistent with $\pi$ such that $s_{0}=s$. Observe that, by definition of the antagonistic and cooperative values, we have:
$\mathbf{aVal}^{\pi}(s)\leq\sum^{\infty}_{t=0}\gamma^{t}R(s_{t},a_{t})\leq\mathbf{cVal}^{\pi}(s).$
The inequalities above hold surely and therefore also hold almost surely. Since the expected value of a constant is that same constant, by the above arguments, we get:
$\mathbf{aVal}^{\pi}(s)=\mathbb{E}[\mathbf{aVal}^{\pi}(s)]\leq\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t},a_{t})\right]\leq\mathbb{E}[\mathbf{cVal}^{\pi}(s)]=\mathbf{cVal}^{\pi}(s).$
Finally, since $\pi$ was arbitrary, then by definition of supremum, we obtain:
$\sup_{\pi}\mathbf{aVal}^{\pi}(s)\leq\sup_{\pi}\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t},a_{t})\right]\leq\sup_{\pi}\mathbf{cVal}^{\pi}(s),$
as required.∎
Appendix B Proof of Theorem 1
Consider a pMDP $\langle S,\iota,A,P,X,R,\gamma\rangle$. Let $\theta\in\Theta_{\mathsf{gp}}$. We will prove that there is a policy $\pi$ such that
$\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})>\mathbf{aVal}(\iota)\right)=1,$
if and only if there is a policy $\pi$ such that $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\right)=1$.
B.1 Preliminaries
We begin with a lemma that will greatly simplify our argument to establish the claim. The first part should be intuitive to the reader since it just says that every action played before (eventually) hitting $\mathcal{I}$ does not decrease the antagonistic value. The second part is less trivial: When a transition from $\mathcal{I}$ is traversed, the (local) antagonistic value may actually increase. Hence, a worst-case optimal policy $\pi$ may choose to play sub-optimally (locally) while still ensuring the best antagonistic value from $\iota$. Thankfully, the intuitive property that $\pi$ can always choose to play optimally from every reached state and not just the initial one does hold true.
Lemma 3.
Let $\pi$ be a policy with $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$. Consider a trajectory $s_{0}a_{0}s_{1}\dots$ consistent with $\pi$ and write $\tau_{\mathcal{I}}\in\mathbb{N}\in\{\infty\}$ for its first hit time to $\mathcal{I}$. Then,
$\forall T<\tau_{\mathcal{I}}:\sum_{t=T}^{\infty}\gamma^{t}R(s_{t},a_{t})=\mathbf{aVal}^{\pi}(s_{T})=\mathbf{aVal}(s_{T}).$ (15)
Moreover, there exists a second policy $\mu$ with $\mathbf{aVal}^{\mu}(\iota)=\mathbf{aVal}(\iota)$, $\operatorname{\mathrm{Pr}}_{\theta}^{\mu}(\tau_{\mathcal{I}}<\infty)=\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(\tau_{\mathcal{I}}<\infty)$, and such that for all trajectories $s_{0}a_{0}s_{1}\dots$ consistent with it have:
$\forall T\in\mathbb{N}:\sum_{t=T}^{\infty}\gamma^{t}R(s_{t},a_{t})\geq\mathbf{aVal}^{\mu}(s_{T})=\mathbf{aVal}(s_{T}),$ (16)
with the inequality being strict if and only if $\langle s_{t},a_{t},s_{t+1}\rangle\in\mathcal{I}$.
Proof.
The first part of the result follows from induction on the definition of $\mathbf{aVal}^{\pi}$, i.e.,
$\mathbf{aVal}^{\pi}(s)=\inf_{\rho}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t}),$
where $\rho$ ranges over trajectories consistent with $\pi$, Equation 8, and Equation 10.
For the second part, we modify $\pi$ to play in a worst-case optimal fashion even after the first transition from $\mathcal{I}$ is traversed. That is, we change action choices so that $\mathbf{aVal}^{\mu}(s)=\mathbf{aVal}(s)$ holds for all $s\in S$. This is an easy local change that is always possible since $\mathbf{aVal}$ is the unique solution to Equation 8 (cf. [49, Sec. 5]). Note that it also preserves the probability of hitting $\mathcal{I}$ since the change is locally applied to actions chosen from states that are only ever reached from $\iota$ via transitions from $\mathcal{I}$ (otherwise, the first part of the claim would make it so that no change is actually needed). Finally, the inequalities then follow from the definition of $\mathcal{I}$ and again an induction on Equation 8 and Equation 10. ∎
B.2 Proof of the claim
We begin with the reverse implication. That is, suppose there is a policy $\pi$ with $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$ such that $\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(\tau_{\mathcal{I}}<\infty)=1$. Without loss of generality, we will also assume our policy enjoys the properties from $\mu$ in Lemma 3. Now, consider an arbitrary trajectory $\rho=s_{0}a_{0}s_{1}\dots$ consistent with $\pi$ and such that it eventually traverses a transition from $\mathcal{I}$. Write $\tau_{\mathcal{I}}$ for the first hit time of $\mathcal{I}$. The following hold:
$\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})={}$ $\sum_{t=0}^{\tau_{\mathcal{I}}-1}\gamma^{t}R(s_{t},a_{t})+\sum_{t=\tau_{\mathcal{I}}}^{\infty}\gamma^{t}R(s_{t},a_{t})$ (17)
[…]
To conclude, write $A$ for the event corresponding to having a trajectory with a value strictly larger than $\mathbf{aVal}(\iota)$ and $B$ for the event of having a finite hit time to $\mathcal{I}$. Recall that we assumed $\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(B)=1$ and we just established that $\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(A\!\mid\!B)=1$. Hence,
$\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(A)\geq\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(A\cap B)=\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(B)\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(A\!\mid\!B)=1,$
and we get the desired result.
To prove the forward implication, we first argue that $\pi$ must be worst-case optimal. That is, we claim that $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$. Towards a contradiction, assume there is a trajectory $\rho=s_{0}a_{0}s_{1}\dots$ consistent with $\pi$ such that $s_{0}=\iota$ and $\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})<\mathbf{aVal}(s_{0})$. The trajectory can be chosen so that there is some $T\in\mathbb{N}$ such that $\mathbf{aVal}(s_{T})-\sum_{t=0}^{T-1}\gamma^{t}R(s_{t},a_{t})>\gamma^{T}\frac{1}{1-\gamma}$. By Equation 10, the worst case is actually witnessed by trajectories composed of a simple path and an infinite repetition of a simple cycle. Hence, after the first divergence from the antagonistic value, discounting ensures such a $T$ indeed exists. Consider the set of all trajectories consistent with $\pi$ and such they have $s_{0}a_{0}s_{1}\dots a_{T-1}s_{T}$ as a prefix. Note that the measure of this event, which we denote $\rho_{T}$ is nonzero, i.e.,
$\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(\rho_{T})=\prod_{t=0}^{T-1}P_{\theta}(s_{t},a_{t},s_{t+1}).$
The fact that it is nonzero follows from our use of graph-preserving valuations only. Now, from our choice of trajectory and $T$, we get that all trajectories in $\rho_{T}$ are such that their discounted sum of rewards is strictly smaller than $\mathbf{aVal}(s_{0})$. We get a contradiction with our assumption that almost all trajectories have a discounted sum of rewards that is, instead, strictly larger.
We have proved that $\pi$ is worst-case optimal. It remains to argue that $\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(\tau_{\mathcal{I}}<\infty)=1$. Towards a contradiction, suppose some nonzero measure of trajectories never hits $\mathcal{I}$. By Lemma 3, Equation 15, we get that the discounted sum of rewards of such trajectories is exactly $\mathbf{aVal}(\iota)$. However, this means we get the same contradiction as above: we have a nonzero measure of trajectories whose discounted sum of rewards is not strictly larger than $\mathbf{aVal}(\iota)$, which goes against our initial assumption.∎
Appendix C Proof of Lemma 2
Suppose $\mathbf{aVal}^{\pi}(\iota)=\mathbf{aVal}(\iota)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\right)=1$ for some policy $\pi$, and recall that the concrete choice of $\theta$ is not relevant for the latter. Now, from the reverse implication of Theorem 1 we get that there is some policy $\pi$ such that:
$\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})>\mathbf{aVal}(\iota)\right)=1.$
It follows that:
$V_{\theta}^{\pi}(\iota)=\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t},a_{t})\right]>\mathbf{aVal}(\iota),$
whence $\mathbf{aVal}(\iota)<V_{\theta}^{\pi}(\iota)\leq V_{\theta}(\iota)$ as required.∎
Appendix D Algorithmically checking Lemma 2
Recall the proposed algorithm is a two-step process:
  1. 1.
    Remove all state-action pairs $\langle s,a\rangle$ for which we have $\mathbf{aVal}(s)>\min_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{aVal}(s^{\prime})$;
  2. 2.
    In the resulting pMDP, use Algorithm 45 for almost-sure reachability from [2].
Termination of the algorithm is trivial since Algorithm 45 also terminates. It remains for us to establish correctness. We proceed by establishing partial correctness and completeness.
D.1 Partial correctness
We start from the correctness guarantee of Algorithm 45 to determine whether there is a policy that ensures almost sure hitting of $\mathcal{I}$ and then establish that the same policy is also worst-case optimal. Let $\pi$ be a policy such that $\operatorname{\mathrm{Pr}}_{\theta}^{\pi}(\tau_{\mathcal{I}}<\infty)=1$ in the sub-pMDP as found by Algorithm 45.
Note that the first step of the procedure can be seen as a restriction of the actions allowed from each state of the pMDP. We formalize this by writing $B(s)\subseteq A$ for the set of actions allowed from state $s\in S$. Now, consider any further restriction $A^{\prime}$ such that $\emptyset\neq A^{\prime}(s)\subseteq B(s)$ for all $s\in S$. In particular, $\mathrm{supp}\circ\pi$ is one such restriction of $B$. From our choice of actions to be removed (resulting in $B$), we have the following.
$\mathbf{aVal}(s)={}$ $\max_{a\in A}\min_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{aVal}(s^{\prime})$ this is just Equation 8 (20)
[…]
It follows that the antagonistic values are also the (unique) solution to the system consisting of Equation 22 for all $s\in S$. This last system is the same as the one prescribed by Equation 10 to obtain the antagonistic value of $\pi$. We thus get that $\mathbf{aVal}(s)=\mathbf{aVal}^{\pi}(s)$ for all $s\in S$, which concludes the proof.∎
D.2 Completeness
This part of the argument is the easiest. Note that the first step of the procedure removes state-action pairs $\langle s,a\rangle$ that would decrease the antagonistic value of the $s$. By Lemma 3, all worst-case optimal policies $\pi$ avoid such state-action pairs, lest some trajectory consistent with them would violate the guarantees of the lemma. This means the set of worst-case optimal policies $\underline{\pi}$ of the pMDP is preserved after removing those state-action pairs. Hence, the procedure does not miss any worst-case optimal policies that hit $\mathcal{I}$ almost surely.∎
Appendix E Proof of Theorem 4
Let $\pi$ be a policy, $\theta\in\Theta_{\mathrm{gp}}$, and $\langle s,a\rangle$ a state-action pair. We prove that if $a\in\mathrm{supp}(\pi(s))$ and $V_{\theta}(s)>Q_{\theta}(s,a),$ then $V_{\theta}(s)>V_{\theta}^{\pi}(s)$.
First, note that $V_{\theta}(s)\geq Q_{\theta}(s,b)\geq Q_{\theta}^{\pi}(s,b)$, for all $b\in A$, by definition of both values. Now, we have the following inequalities:
$V^{\pi}_{\theta}(s)={}$ $\sum_{b\in A}\pi(b\!\mid\!s)Q^{\pi}_{\theta}(s,b)$ by definition
[…]
This concludes the proof. ∎
Appendix F Proof of Theorem 3
Let $\pi$ be a policy, $\theta\in\Theta_{\mathrm{gp}}$, and $\langle s,a\rangle$ a state-action pair. We prove that if $a\in\mathrm{supp}(\pi(s))$ and
$\mathbf{aVal}(s)>Q_{\theta}(s,a),$
then $V_{\theta}(s)>V_{\theta}^{\pi}(s)$. Moreover, if $\mathbf{aVal}^{\pi}(s)=\mathbf{aVal}(s)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\!\mid\!s_{0}=s\right)=1$, for some $\pi$, the nonstrict version of the equation suffices.
From Lemma 1 and the condition of the claim, it follows that $V_{\theta}(s)>Q_{\theta}(s,a)$. Then, Theorem 4 implies the result. For the second part of the claim, we make use of Lemma 2 to get $V_{\theta}(s)>\mathbf{aVal}(s)\geq Q_{\theta}(s,a)$ before again invoking Theorem 4. ∎
Appendix G Proof of Theorem 2
Let $\pi$ be a policy and $\langle s,a\rangle$ be a state-action pair. We prove that if $a\in\mathrm{supp}(\pi(s))$ and
$\mathbf{aVal}(s)>\max_{s^{\prime}\in\mathrm{supp(s,a)}}R(s,a)+\gamma\mathbf{cVal}(s^{\prime}),$
then for all $\theta\in\Theta_{\mathrm{gp}}$ we have $V_{\theta}(s)>V_{\theta}^{\pi}(s)$. Moreover, if $\mathbf{aVal}^{\pi}(s)=\mathbf{aVal}(s)$ and $\operatorname{\mathrm{Pr}}^{\pi}_{\theta}\left(\tau_{\mathcal{I}}<\infty\!\mid\!s_{0}=s\right)=1$, for some $\pi$, the nonstrict version of the equation suffices.
Assume that the condition of the claim is true. We get the following relations.
$Q_{\theta}(s,a)={}$ $R(s,a)+\gamma\sum_{s^{\prime}\in S}P_{\theta}(s^{\prime}\!\mid\!s,a)V_{\theta}(s^{\prime})$ by definition
[…]
We have just established that $\mathbf{aVal}(s)>Q_{\theta}(s,a)$ for the first part of the claim and that the nonstrict version of the inequality holds for the second part of the claim. The result thus follows from Theorem 3.∎
Appendix H Overview of the benchmarks
Behavior policies.
For each environment, we construct a behavior policy to collect the datasets and serve as a baseline for (p)SPIBB to bootstrap. This behavior policy is obtained by first computing an optimal (deterministic) policy using policy iteration. For each state, we make the policy stochastic by perturbating the probability mass by some parameter $\alpha\in[0,1]$. Specifically, we remove probability mass $\alpha$ from the optimal action and re-distribute it uniformly over the other actions. Formally, the behavior policy is constructed as follows:
$\pi_{B}(a\mid s)=\begin{cases}1-\alpha&\textbf{if }\pi(a\mid s)=1,\\ \frac{\alpha}{\lvert A\rvert}&\textbf{if }\pi(a\mid s)=0.\end{cases}$
An overview of the dimensions and perturbation parameter $\alpha$ of all benchmarks can be found in Table 1. All benchmarks, except Rock-Paper-Scissors, were derived from existing literature, as referenced in the main text.
Benchmark Source $\lvert S\rvert$ $\lvert A\rvert$ $\lvert X\rvert$ $\alpha$
Gridworld [24] 25 4 1 $\nicefrac{{1}}{{2}}$
Resource gathering [6] 376 4 1 $\nicefrac{{1}}{{5}}$
Pac-Man [19] 498 5 0 $\nicefrac{{1}}{{20}}$
Rock-paper-scissors This paper 1321 3 9 $\nicefrac{{1}}{{20}}$
Taxi [16] 501 6 300 $\nicefrac{{1}}{{20}}$
Table 1: Overview of the different benchmarks.
Rock-Paper-Scissors.
Rock-Paper-Scissors (RPS) is a custom benchmark we introduce in this paper. The RPS benchmark is constructed as a $20$-round game of rock-paper-scissors, in which the player receives a reward of $+1$ for winning, $-1$ for losing, and 00 in case of a draw. The opponent of the player stochastically chooses whether to play rock, paper, or scissors, making use of a non-uniform distribution. This non-uniformity induces a bias that the player can exploit. This experiment of finding an optimal strategy while playing against a biased opponent is inspired by a paper [9], where the authors describe experiments where human participants had to play RPS against a biased bot. The bias used in the RPS benchmarks is similar to an experiment in that paper and involves the opponent being more likely to choose a play that would have beaten the player in the previous round.